# Deliberate Planning of 3D Bin Packing on Packing Configuration Trees

### v3

@@ Abstract
Online 3D Bin Packing Problem (3D-BPP) has widespread applications in industrial automation and has aroused enthusiastic research interest recently. Existing methods usually solve the problem with limited resolution of spatial discretization, and/or cannot deal with complex practical constraints well. We propose to enhance the practical applicability of online 3D-BPP via learning on a novel hierarchical representation—packing configuration tree (PCT). PCT is a full-fledged description of the state and action space of bin packing which can support packing policy learning based on deep reinforcement learning (DRL). The size of the packing action space is proportional to the number of leaf nodes, i.e., candidate placements, making the DRL model easy to train and well-performing even with continuous solution space. We further discover the potential of PCT as tree-based planners in deliberately solving packing problems of industrial significance, including large-scale packing and different variations of BPP setting. A recursive packing method is proposed to decompose large-scale packing into smaller sub-trees while a spatial ensemble mechanism integrates local solutions into a global one. For different BPP variations with additional decision variables, such as lookahead, buffering, and offline packing, we propose a unified planning framework enabling out-of-the-box problem solving based on a pre-trained PCT model. Extensive evaluations demonstrate that our method outperforms existing online BPP baselines and is versatile in incorporating various practical constraints. Driven by PCT, the planning process excels across large-scale problems and diverse problem variations, with performance improving as the problem scales up and the decision variables grow. To verify our method, we develop a real-world packing robot for industrial warehousing, with careful designs accounting for constrained placement and transportation stability. Our packing robot operates reliably and efficiently on unprotected pallets at 10 seconds per box. It achieves averagely 19 boxes per pallet with 57.4% space utilization for relatively large-size boxes.
@@ 1 Introduction
As one of the most classic combinatorial optimization problems, the 3D bin packing problem usually refers to packing a set of cuboid-shaped items $i\in\mathcal{I}$, with sizes $s^{x}_{i},s^{y}_{i},s^{z}_{i}$ along $x,y,z$ axes, respectively, into the maximum space utilization of bin $C$ with sizes $S^{x},S^{y},S^{z}$, in an axis-aligned fashion. Traditional 3D-BPP assumes that all the items to be packed are known a priori (Martello et al., 2000), which is also called offline BPP. The problem is known to be strongly NP-hard  (De Castro Silva et al., 2003). However, in many real-world application scenarios, e.g., logistics or warehousing (Wang and Hauser, 2019a), the upcoming items cannot be fully observed; only the current item to be packed is observable. Packing items without the knowledge of all upcoming items is referred to as online BPP (Seiden, 2002).
@@
We propose to enhance learning-based online 3D-BPP towards practical applicability through learning with a novel hierarchical representation––packing configuration tree (PCT). PCT is a dynamically growing tree where the internal nodes describe the space configurations of packed items and leaf nodes the packable placements of the current item. PCT is a full-fledged description of the state and action space of bin packing which can support packing policy learning based on deep reinforcement learning (DRL). We extract state features from PCT using graph attention networks (Velickovic et al., 2018) which encode the spatial relations of all configuration nodes. The state feature is input into the actor and critic networks of the DRL model. The actor network, designed based on a pointer mechanism, weighs the leaf nodes and outputs the final placement.
@@
PCT was published in ICLR 2022 (Zhao et al., 2022a), which is the first learning-based method that successfully solves online 3D-BPP with continuous solution space and achieves strong performance. We believe that its potential extends beyond regular online packing and further discover its capability as tree-based planners to deliberately solve packing problems of industrial significance, including large-scale packing and different variations of BPP setting. We propose recursive packing to decompose the tree structure of large-scale online 3D-BPP as smaller sub-trees to individually solve them. The obtained local solutions are then integrated into a global one through an effective spatial ensemble method. In addition to online packing, PCT’s enhanced representation of packing constraints and flexible action space can be extended to other mainstream BPP settings, such as lookahead, buffering, and offline packing. We propose a unified planning framework enabling out-of-the-box problem solving based on a pre-trained PCT model.
We have established a real-world packing robot in an industrial warehouse, carefully designed to meet constrained placement (Choset et al., 2005) and transportation stability (Hof et al., 2005) requirements. Unlike laboratory setups with protective container walls (Yang et al., 2021a; Xu et al., 2023), our system operates under industrial standards with boxes (items) directly placed onto unprotected pallets. Even minor robot-object collisions during placement can destabilize the static stack, and dynamic transportation by Automated Guided Vehicles (AGVs) or human workers may further challenge the stack’s stability. To satisfy constrained placement, our system incorporates a modular end-effector capable of actively adjusting its shape to maximize gripping force while minimizing collision risk. To ensure transportation stability, we perform physics-based verification via test-time simulation to account for real-world uncertainties. Each placement is evaluated under multiple sets of disturbances, with the simulation accelerated by GPU-based batch parallelism (Makoviychuk et al., 2021). Combined with an asynchronous decision-making pipeline that overlaps decision time with robot execution, our packing robot operates efficiently and reliably on unprotected pallets in industrial settings, with a cycle time of 10 seconds per box and averagely 19 boxes per pallet (57.4% space utilization for relatively large-size boxes). A video showcasing our dynamic packing process is provided at this URL.
@@
    We propose a unified planning framework to solve different BPP variations out of the box, based on a pre-trained PCT model without fine-tuning.
@@ 2 Related Work
Given a single bin $C$ and a set of items $\mathcal{I}$, the objective of the 3D bin packing problem (3D-BPP) (Martello et al., 2000) is to maximize the space utilization. Its basic constraints can be formulated as follows:
@@
where $p_{i}$ denotes the Front-Left-Bottom (FLB) coordinate of item $i\in\mathcal{I}$, and $N$ is the total number of items after packing. The variable $d\in\{x,y,z\}$ represents the axis. If item $i$ is placed before item $j$ along axis $d$, the value of $e_{ij}^{d}$ is 1; otherwise, it is 0. Equations 2 and 3 represent the non-overlapping constraint and containment constraint (Martello et al., 2000), respectively.
@@ 2.2 Learning-based Online Packing
The heuristic methods are intuitive to implement and can be easily applied to various scenarios. However, the price of good flexibility is that these methods perform mediocrely, especially for online 3D-BPP with specific constraints. Designing new heuristics for specific classes of 3D-BPP is heavy work since this problem has an NP-hard solution space where many situations need to be premeditated manually by trial and error. Substantial domain knowledge is also necessary to ensure safety and reliability. To automatically generate a policy that works well on specified online 3D-BPP, Verma et al. (2020) and Zhao et al. (2021) employ DRL to solve this problem, however, their methods only work in discrete and small coordinate spaces. Despite their limitations, these works are soon followed for logistics robot implementation (Hong et al., 2020; Yang et al., 2021b; Zhao et al., 2022b). Referring to Hu et al. (2017), Zhang et al. (2021) adopt a online placement policy for offline packing needs. All these learning-based methods work in a grid world with limited discretization accuracy, which reduces their practical applicability. PCT overcomes these limitations by explicitly storing the necessary packing configuration information in a structured packing representation and using graph neural networks to capture spatial relationships. This allows PCT to better represent the packing state, enhancing the performance of DRL.
PCT is the first learning-based method to successfully solve online 3D-BPP in continuous solution space. Its core idea is to identify a finite set of candidates from the continuous domain and use DRL to determine the best solution. This candidate-based packing mechanism has been widely adopted in subsequent work. Yuan et al. (2023) and Pan et al. (2023b) optimize PCT policies from specific perspectives about performance variance and lower bounds. TAP-NET++ (Xu et al., 2023) extends this approach by simultaneously calculating the attention between multiple candidates and multiple items to address buffering packing (Puche and Lee, 2022). Zhao et al. (2023) theoretically prove the local optimality of limited candidates for packing irregularly shaped items, using DRL to find the global solution. SDF-Pack (Pan et al., 2023a) focuses on finding a currently greedy, compact solution, but the most compact placement at a given moment is not necessarily optimal for the entire packing sequence.
@@ 2.3 Practical Constraints of Industrial Packing
The majority of literature on 3D-BPP (Martello et al., 2000) focuses primarily on basic non-overlapping constraint 2 and containment constraint 3. Failing to consider essential real-world constraints, such as stability (Ramos et al., 2016), these algorithms have limited industrial applicability. Zhao et al. (2022b) propose a fast quasi-static equilibrium estimation method tailored for DRL training and test their learned policies with real logistics boxes. A key limitation of their approach is the use of a heightmap (the upper frontier of packed items) state representation, similar to Zhang et al. (2021), which overlooks the underlying constraints between packed items. The lack of spatial information in this representation makes the problem a partially observable Markov Decision Process (POMDP) (Spaan, 2012), which complicates DRL training and limits performance on more complex practical 3D-BPP instances involving constraints like isle friendliness and load balancing (Gzara et al., 2020). PCT overcomes these limitations by explicitly storing the necessary packing configuration information in a tree structure and using graph attention networks (Velickovic et al., 2018) to capture spatial relationships. This allows PCT to better represent the packing state, thereby enhancing the performance of DRL.
@@
Large-scale packing plays a critical role in production. For truck packing (Egeblad et al., 2007), hundreds of items must be packed online before long-distance transportation. However, DRL methods face challenges when applied to large-scale combinatorial optimization (CO) tasks (Kool et al., 2019; Qiu et al., 2022). On the one hand, exploring via trial and error struggles to collect sufficient learning samples in the enormous NP-hard space. On the other hand, long sequences of decision-making leads to learning instability (Sutton and Barto, 2018). While recent studies demonstrate that graph-based neural solvers exhibit problem scale generalizability (Sun and Yang, 2023), performance degradation is still observed due to test distribution mismatches (Yu et al., 2024). Leveraging the structured packing representation of PCT, we recursively decompose large-scale packing into smaller sub-trees and integrate local solutions into a global one using a spatial ensemble mechanism. This deliberate planning achieves state-of-the-art performance on large-scale packing, with performance continuing to improve as the the problem scales up.
@@
In industrial applications, strictly online BPP (Seiden, 2002) is not the only demand, and additional packing settings along with decision variables need to be considered. For example, lookahead packing (Grove, 1995) allows for observing upcoming items in advance for enabling better placement of the current one. Buffereing packing (Puche and Lee, 2022) temporarily stores incoming items in a buffer, allowing the robot to select any one of them within reach. Offline packing (Martello et al., 2000; Demisse et al., 2012) receives complete item information from the central control system to schedule their arrival order. While various solvers have been proposed recently to address these settings, they typically rely on additional parameterized modules (Hu et al., 2017; Duan et al., 2019) to handle the extra constraints and decision variables, increasing training overhead and decreasing transferability. The advantages of PCT—better constraint representation and more flexible decision-making—can also benefit these settings. We induce these problems into distinct planning processes, where the constraints do not conflicts, allowing problems solved with a unified framework out of the box. Without fine-tuning requirement, this framework achieves consistency state-of-the-art performance across various BPP settings.
@@ 3 Method
[…]
We begin by introducing our learning-based packing solver, PCT, and demonstrating its enhanced representation of packing constraints and flexible action space in Section 3.1. In Section 3.2, we formalize PCT-based packing as a Markov Decision Process (MDP) for policy learning. Using pre-trained PCT models, we then perform deliberate planning to solve packing problems with industrial significance: large-scale packing and different variations of BPP setting. In Section 3.3, we present a recursive packing method to tackle large-scale challenges, along with a spatial ensemble mechanism to integrate local solutions into a global one. In Section 3.4, we introduce a unified planning framework that solves different BPP variations out of the box.
@@ 3.1 Packing Configuration Tree
When a rectangular item $n_{t}$ is added to a given packing with position $(p^{x}_{n},p^{y}_{n},p^{z}_{n})$ at time step $t$, it introduces a series of new candidate positions where future items can be accommodated, as illustrated in Figure 1. Combined with the axis-aligned orientation $o\in\textbf{O}$ for $n_{t}$ based on existing positions, we get candidate placements (i.e. position and orientation). The packing process can be seen as a placement node being replaced by a packed item node, and new candidate placement nodes are generated as children. As the packing time step $t$ goes on, these nodes are iteratively updated and a dynamic packing configuration tree is formed, denoted as $\mathcal{T}$. The internal node set $\textbf{B}_{t}\in\mathcal{T}_{t}$ represents the space configurations of packed items, and the leaf node set $\textbf{L}_{t}\in\mathcal{T}_{t}$ the packable candidate placements. During the packing, leaf nodes that are no longer feasible, e.g., covered by packed items, will be removed from $\textbf{L}_{t}$. When there is no packable leaf node that makes $n_{t}$ satisfy the constraints of placement, the packing episode ends. Without loss of generality, we stipulate a vertical top-down packing within a single bin (Wang and Hauser, 2021).
@@
The performance of online 3D-BPP policies has a strong relationship with the choice of leaf node expansion schemes––which incrementally calculate new candidate placements introduced by the just placed item $n_{t}$. A good expansion scheme should reduce the number of solutions to be explored while not missing too many feasible packings. Meanwhile, polynomials computability is also expected. Designing such a scheme from scratch is non-trivial. Fortunately, several placement rules independent from particular packing problems have been proposed, such as Corner Point  (Martello et al., 2000), Extreme Point (Crainic et al., 2008), and Empty Maximal Space (Ha et al., 2017). We extend these schemes, which have proven to be accurate and efficient, to our PCT expansion. Their performance will be reported in Section 4.1.
@@
Given the bin configuration $\mathcal{T}_{t}$ and the current item $n_{t}$, the packing policy can be parameterized as $\pi(\textbf{L}_{t}|\mathcal{T}_{t},n_{t})$. The tuple $(\mathcal{T}_{t},n_{t})$ can be treated as a graph and encoded by Graph Neural Networks (GNNs) (Gori et al., 2005). Specifically, the PCT keeps growing with time step $t$ and cannot be embedded by spectral-based approaches (Bruna et al., 2014) requiring fixed graph structure. We adopt non-spectral Graph Attention Networks (GATs) (Velickovic et al., 2018), which require no graph structure priori.
@@
$\textbf{h}^{\prime}=\hat{\textbf{h}}+\text{GAT}(\hat{\textbf{h}}),\quad\textbf{h}=\textbf{h}^{\prime}+\phi_{FF}(\textbf{h}^{\prime}),\vspace{-4pt}$ (5)
where $\phi_{FF}$ is a node-wise Feed-Forward MLP with output dimension $d_{h}$ and $\textbf{h}^{\prime}$ is an intermediate variable. Equation 5 can be seen as an independent block and be repeated multiple times with different parameters. We don’t extend GAT to employ the multi-head attention mechanism (Vaswani et al., 2017) since we find that additional attention heads cannot help the final performance. We execute Equation 5 once and we set $d_{v}=d_{k}$. More implementation details are provided in Appendix A.
@@
$q=W^{q}\bar{h},\quad k_{i}=W^{k}h_{i},\quad u_{i}=\frac{q^{T}k_{i}}{\sqrt{d_{\text{k}}}}.\vspace{-6pt}$ (6)
Here $h_{i}$ only comes from $\textbf{h}_{\textbf{L}}$. The compatibility vector $\textbf{u}_{\textbf{L}}\in\mathbb{R}^{|\textbf{L}_{t}|}$ represents the leaf node selection logits. The probability distribution over the PCT leaf nodes $\textbf{L}_{t}$ is:
$\pi_{\theta}(\textbf{L}_{t}|\mathcal{T}_{t},n_{t})=softmax\left(c_{clip}\cdot\tanh\left(u_{\textbf{L}}\right)\right).\vspace{-4pt}$ (7)
Following Bello et al. (2017), the compatibility logits are clipped with tanh, where the range is controlled by hyperparameter $c_{clip}$, and finally normalized by softmax.

### v4

@@ Abstract
The online 3D Bin Packing Problem (3D-BPP) has widespread applications in industrial automation and has aroused enthusiastic research interest recently. Existing methods usually solve the problem with limited resolution of spatial discretization, and/or cannot deal with complex practical constraints well. We propose to enhance the practical applicability of online 3D-BPP via learning on a novel hierarchical representation—packing configuration tree (PCT). PCT is a full-fledged description of the state and action space of bin packing which can support packing policy learning based on deep reinforcement learning (DRL). The size of the packing action space is proportional to the number of leaf nodes, i.e., candidate placements, making the DRL model easy to train and well-performing even with continuous solution space. We further discover the potential of PCT as tree-based planners in deliberately solving packing problems of industrial significance, including large-scale packing and different variations of the BPP setting. A recursive packing method is proposed to decompose large-scale packing into regular sub-trees while a spatial ensemble mechanism integrates local solutions into a global one. For different BPP variations with additional decision variables, such as lookahead, buffering, and offline packing, we propose a unified planning framework enabling various problem solving based on a pre-trained PCT model with no additional adaptation. Extensive evaluations demonstrate that our method outperforms existing online BPP baselines and is versatile in incorporating various practical constraints. Driven by PCT, the planning process excels across large-scale problems and diverse problem variations, with performance improving as the problem scales up and the decision variables grow. To verify our method, we develop a real-world packing robot for industrial warehousing, with careful designs accounting for constrained placement and transportation stability. Our packing robot operates reliably and efficiently on unprotected pallets at 9.8 seconds per box. It achieves averagely 19 boxes per pallet with 57.4% space utilization for large-size boxes.
@@ 1 Introduction
As one of the most classic combinatorial optimization problems, the 3D bin packing problem (3D-BPP) usually refers to packing a set of cuboid-shaped items $i\in\mathcal{I}$, with sizes $s^{x}_{i},s^{y}_{i},s^{z}_{i}$ along $x,y,z$ axes, respectively, into the maximum space utilization of bin $C$ with sizes $S^{x},S^{y},S^{z}$, in an axis-aligned fashion. Traditional 3D-BPP assumes that all the items to be packed are known a priori (Martello et al., 2000), which is also called offline BPP. The problem is known to be strongly NP-hard  (De Castro Silva et al., 2003). However, in many real-world application scenarios, e.g., logistics or warehousing (Wang and Hauser, 2019a), the upcoming items cannot be fully observed; only the current item to be packed is observable. Packing items without the knowledge of all upcoming items is referred to as online BPP (Seiden, 2002).
@@
We propose to enhance learning-based online 3D-BPP towards practical applicability through learning with a novel hierarchical representation–packing configuration tree (PCT). PCT is a dynamically growing tree where the internal nodes describe the space configurations of packed items and leaf nodes the packable placements of the current item. It is a full-fledged description of the state and action space of bin packing which can support packing policy learning based on deep reinforcement learning (DRL). We extract state features from PCT using graph attention networks (Velickovic et al., 2018) which encode the spatial relations of all configuration nodes. The state feature is input into the actor and critic networks of the DRL model. The actor network, designed based on a pointer mechanism, weighs the leaf nodes and outputs the final placement.
@@
PCT was published in ICLR 2022 (Zhao et al., 2022a), which is the first learning-based method that successfully solves online 3D-BPP with continuous solution space and achieves strong performance. We believe that its potential extends beyond regular online packing and further discover its capability as tree-based planners to deliberately solve packing problems of industrial significance, including large-scale packing and different variations of BPP setting. We propose recursive packing to decompose the tree structure of large-scale online 3D-BPP as regular sub-trees to individually solve them. The obtained local solutions are then integrated into a global one through an effective spatial ensemble method. In addition to online packing, PCT’s enhanced representation of packing constraints and flexible action space can be extended to other mainstream BPP settings, such as lookahead, buffering, and offline packing. We propose a unified planning framework enabling various problem solving based on a pre-trained PCT model with no additional adaptation.
We establish a real-world packing robot in an industrial warehouse, carefully designed to meet constrained placement (Choset et al., 2005) and transportation stability (Hof et al., 2005) requirements. Unlike laboratory setups with protective container walls (Yang et al., 2021b; Xu et al., 2023), our system operates under industrial standards with boxes (items) directly placed onto unprotected pallets. Even minor robot-object collisions during placement can destabilize the static stack, and dynamic transportation by Automated Guided Vehicles (AGVs) or human workers may further challenge the stack’s stability. To satisfy constrained placement, our system incorporates a modular end-effector capable of actively adjusting its shape to maximize gripping force while minimizing collision risk. To ensure transportation stability, we perform physics-based verification via test-time simulation to account for real-world uncertainties. Each placement is evaluated under multiple sets of disturbances, with the simulation accelerated by GPU-based batch parallelism (Makoviychuk et al., 2021). Combined with an asynchronous decision-making pipeline that overlaps decision time with robot execution, our packing robot operates efficiently and reliably on unprotected pallets in industrial settings, with a cycle time of 9.8 seconds per box and averagely 19 boxes per pallet (57.4% space utilization for large boxes). A dynamic packing video is provided at URL.
@@
    We propose a unified planning framework to solve different BPP variations, based on a pre-trained PCT model with no additional adaptation.
@@ 2 Related Work
Given a single bin $C$ and a set of items $\mathcal{I}$, the objective of 3D-BPP (Martello et al., 2000) is to maximize the space utilization. Its basic constraints can be formulated as follows:
@@
where $p_{i}$ denotes the Front-Left-Bottom (FLB) coordinate of item $i\in\mathcal{I}$, and $N$ is the total number of items after packing. The variable $d\in\{x,y,z\}$ represents the axis. If item $i$ is placed before item $j$ along axis $d$, the value of $e_{ij}^{d}$ is 1; otherwise, it is 0. Equations (2) and (3) represent the non-overlapping constraint and containment constraint (Martello et al., 2000), respectively.
@@ 2.2 Learning-based Online Packing
The heuristic methods are intuitive to implement and can be easily applied to various online 3D-BPP scenarios. However, the price of good flexibility is that these methods perform poorly since they cannot be adapted to specific problem distributions. In later sections, we show that even the best-performing heuristics underperform our learning-based PCT method by at least 11.5% across all online 3D-BPP test instances. Designing new heuristics for specific classes of 3D-BPP is heavy work since this problem has an NP-hard solution space where many situations need to be premeditated manually by trial and error. Substantial domain knowledge is also necessary to ensure safety and reliability. To automatically generate a policy that works well on specified online 3D-BPP, Verma et al. (2020) and Zhao et al. (2021) employ DRL to solve this problem; however, their methods only work in discrete and small coordinate spaces. Despite their limitations, these works are soon followed for logistics robot implementation (Hong et al., 2020; Yang et al., 2021b; Zhao et al., 2022b). Referring to Hu et al. (2017), Zhang et al. (2021) adopt an online placement policy for offline packing needs. All these methods work in a grid world with limited discretization accuracy, which reduces their practical applicability.
PCT is the first learning-based method to successfully solve online 3D-BPP in continuous solution space. Its core idea is to identify a finite set of candidates from the continuous domain and use DRL to determine the best solution. This candidate-based packing mechanism has been widely adopted in subsequent work. Yuan et al. (2023) and Pan et al. (2023) optimize PCT policies from specific perspectives about performance variance and the worst performance. TAP-NET++ (Xu et al., 2023) extends this approach by simultaneously calculating the attention between multiple candidates and multiple items to address buffering packing (Puche and Lee, 2022). Zhao et al. (2023) theoretically prove the local optimality of limited candidates for packing irregularly shaped items, using DRL to further determine the global solution.
@@ 2.3 Practical Constraints of Industrial Packing
The majority of literature on 3D-BPP (Martello et al., 2000) focuses primarily on basic non-overlapping constraint (2) and containment constraint (3). Failing to consider essential real-world constraints, such as stability (Ramos et al., 2016), these algorithms have limited industrial applicability. Zhao et al. (2022b) propose a fast quasi-static equilibrium estimation method tailored for DRL training and test their learned policies with real logistics boxes. A key limitation of their approach is the use of a heightmap (the upper frontier of packed items) state representation, similar to Zhang et al. (2021), which overlooks the underlying constraints between packed items. The lack of spatial information makes the problem a partially observable Markov Decision Process (POMDP) (Spaan, 2012), which complicates DRL training and limits performance on more complex practical 3D-BPP instances involving constraints like isle friendliness and load balancing (Gzara et al., 2020). PCT overcomes these limitations by explicitly storing the necessary packing configuration in a tree structure and using graph attention networks (Velickovic et al., 2018) to capture their relationships. This allows PCT to better represent the packing state and enhance the performance of DRL.
@@
Large-scale packing plays a common role in industrial production. For truck packing (Egeblad et al., 2007), hundreds of items must be packed online before long-distance transportation. However, DRL faces challenges when applied to large-scale combinatorial optimization (CO) tasks (Kool et al., 2019; Qiu et al., 2022). On the one hand, learning via trial and error struggles to collect sufficient samples in the enormous NP-hard space. On the other hand, long sequences of decision-making lead to learning instability (Sutton and Barto, 2018). While recent studies demonstrate that graph-based neural solvers exhibit problem scale generalizability (Zhao et al., 2024), performance degradation is still observed due to test distribution mismatches. Leveraging the structured packing representation of PCT, we can decompose the large-scale packing into regular sub-problems to effectively solve them and integrate local solutions into a global one. This deliberate planning achieves state-of-the-art performance on large-scale packing, with performance continuing to improve as the problem scale increases.
@@
In industrial applications, strictly online BPP (Seiden, 2002) is not the only demand. Packing settings along with additional decision variables usually need to be considered. For example, lookahead packing (Grove, 1995) allows for observing upcoming items in advance for enabling better placement of the current one. Buffering packing (Puche and Lee, 2022) temporarily stores incoming items in a buffer, allowing the robot to select any one of them within reach. Offline packing (Martello et al., 2000; Demisse et al., 2012) receives complete item information from the central control system to schedule placement order. While various solvers have been proposed recently to address these settings, they typically rely on additional parameterized modules (Hu et al., 2017; Duan et al., 2019) to handle the extra constraints and decision variables, increasing training overhead and decreasing transferability. The advantages of PCT—better constraint representation and more flexible decision-making—can also benefit these settings. We induce these problems into non-conflicting planning processes within a unified framework, achieving consistently state-of-the-art performance across various BPP settings.
@@ 3 Method
Figure 1: PCT expansion illustrated using a 2D example (in xozxoz plane) for simplicity, and the number of allowed orientations $|\textbf{O}|$ is 1 (see Appendix B for the 3D version). A newly added item introduces a series of empty spaces and new candidate placements are generated, e.g., the left-bottom corner of the empty space.
We begin by introducing our learning-based packing solver, PCT, and demonstrating its enhanced representation of packing constraints and flexible action space in Section 3.1. In Section 3.2, we formalize PCT-based packing as a Markov Decision Process (MDP) for policy learning. Using pre-trained PCT models, we then perform deliberate planning to solve packing problems with industrial significance: large-scale packing and different variations of the BPP setting. In Section 3.3, we present a recursive packing method to tackle large-scale challenges, along with a spatial ensemble mechanism to integrate local solutions into a global one. In Section 3.4, we introduce a unified planning framework that solves different BPP variations with no additional adaptation.
@@ 3.1 Packing Configuration Tree
When a rectangular item $n_{t}$ is added to a given packing with position $(p^{x}_{n},p^{y}_{n},p^{z}_{n})$ at time step $t$, it introduces a series of new candidate positions where future items can be accommodated, as illustrated in Figure 1. Combined with the axis-aligned orientation $o\in\textbf{O}$ for $n_{t}$ based on existing positions, we get candidate placements (i.e., position and orientation). The packing process can be seen as a placement node being replaced by a packed item node, and new candidate placement nodes are generated as children. As the packing time step $t$ goes on, these nodes are iteratively updated and a dynamic packing configuration tree is formed, denoted as $\mathcal{T}$. The internal node set $\textbf{B}_{t}\in\mathcal{T}_{t}$ represents the space configurations of packed items, and the leaf node set $\textbf{L}_{t}\in\mathcal{T}_{t}$ the packable candidate placements. During the packing, leaf nodes that are no longer feasible, e.g., covered by packed items, will be removed from $\textbf{L}_{t}$. When there is no packable leaf node that makes $n_{t}$ satisfy the constraints of placement, the packing episode ends. Without loss of generality, we stipulate a top-down packing within a single bin (Wang and Hauser, 2021).
@@
The performance of online 3D-BPP policies has a strong relationship with the choice of leaf node expansion schemes, which incrementally calculate new candidate placements introduced by the just placed item $n_{t}$. A good expansion scheme should reduce the number of solutions to be explored while not missing too many feasible packings. Meanwhile, polynomials computability is also expected. Designing such a scheme from scratch is non-trivial. Fortunately, several placement rules independent from particular packing problems have been proposed, such as Corner Point  (Martello et al., 2000), Extreme Point (Crainic et al., 2008), and Empty Maximal Space (Ha et al., 2017). We extend these schemes, which have proven to be accurate and efficient, to our PCT expansion. Their performance will be reported in Section 4.1.
@@
Given the bin configuration $\mathcal{T}_{t}$ and the current item $n_{t}$, the packing policy can be parameterized as $\pi(\textbf{L}_{t}|\mathcal{T}_{t},n_{t})$. The tuple $(\mathcal{T}_{t},n_{t})$ can be treated as a graph and encoded by Graph Neural Networks (GNNs) (Gori et al., 2005). Specifically, the PCT keeps growing with time step $t$ and cannot be embedded by spectral-based approaches (Bruna et al., 2014) requiring a fixed graph structure. We adopt non-spectral Graph Attention Networks (GATs) (Velickovic et al., 2018), which require no graph structure priori.
@@
$\textbf{h}^{\prime}=\hat{\textbf{h}}+\text{GAT}(\hat{\textbf{h}}),\quad\textbf{h}=\textbf{h}^{\prime}+\phi_{FF}(\textbf{h}^{\prime}),\vskip-4.0pt$ (5)
where $\phi_{FF}$ is a node-wise Feed-Forward MLP with output dimension $d_{h}$ and $\textbf{h}^{\prime}$ is an intermediate variable. Since each node needs to calculate its relation with all other nodes, the complexity of this GAT operator is $O(N^{2})$. Equation (5) can be seen as an independent block and can be repeated multiple times with different parameters. We don’t extend GAT to employ the multi-head attention mechanism (Vaswani et al., 2017) since we find that additional attention heads cannot help the final performance. We execute Equation (5) once and we set $d_{v}=d_{k}$. More implementation details are provided in Appendix A.
@@
$q=W^{q}\bar{h},\quad k_{i}=W^{k}h_{i},\quad u_{i}=\frac{q^{T}k_{i}}{\sqrt{d_{\text{k}}}}.\vskip-6.0pt$ (6)
Here $h_{i}$ only comes from $\textbf{h}_{\textbf{L}}$. The compatibility vector $\textbf{u}_{\textbf{L}}\in\mathbb{R}^{|\textbf{L}_{t}|}$ represents the leaf node selection logits. This pointer calculation possesses a complexity of $O(|\textbf{L}_{t}|)$. The probability distribution over the PCT leaf nodes $\textbf{L}_{t}$ is:
$\pi_{\theta}(\textbf{L}_{t}|\mathcal{T}_{t},n_{t})=softmax\left(c_{clip}\cdot\tanh\left(u_{\textbf{L}}\right)\right).\vskip-4.0pt$ (7)
Following Bello et al. (2017), the compatibility logits are clipped with tanh, where the range is controlled by hyperparameter $c_{clip}$, and finally normalized by softmax. We provide the computational costs of GAT and the pointer mechanism in Appendix D.